# Lensing and light rings of parity-odd rotating boson stars

### v2

@@ Abstract
We obtain the first image of a parity-odd celestial body. Recently, an intriguing parity-odd rotating boson star has been proposed, and we investigate the lensing effects of these stars in detail. Our analysis demonstrates distinct gravitational distortions around these stars, clearly differentiating them from their parity-even counterparts. Furthermore, we analyze the conditions under which chaotic behavior appears in the images of ultra-compact stars.
@@ II.2 Numerical approach
where $\bar{W}=W/r^{2}$. The parameter space of the parity-odd RBS is shown in Fig.1, which displays a typical inspiralling pattern over the frequency range $w_{\text{min}}<w<1$ with $w_{\text{min}}\approx 0.6835$.
Figure 1: Parity-odd RBS solutions in an Arnowitt-Deser-Misner (ADM) mass $M$ vs scalar field frequency $w$ diagram. Points $1$-$13$ corresponds to the BSs to be discussed, see Table 1 for details.
@@ Acknowledgements.
YH would like to thank Professor Carlos Herdeiro for his prompt and helpful correspondence. This work is supported by National Natural Science Foundation of China Grants Nos.12275106 and 12235019, Shandong Provincial Natural Science Foundation under grant No. ZR2024QA032, as well as Shandong Provincial Natural Science Foundation under grant No. ZR201709220395. DJL is supported by the innovation program of Shanghai Normal University under Grant No. KF202472. Some of the calculations were performed in High-performance Computing Platform at University of Jinan.

### v3

@@ Abstract
We obtain the first image of a parity-odd celestial body. Recently, an intriguing parity-odd rotating boson star was proposed. We investigate the lensing effects of these stars in detail. Our analysis demonstrates distinct gravitational distortions around these stars, clearly differentiating them from their parity-even counterparts. Furthermore, we analyze the conditions under which chaotic behavior appears in the images of ultra-compact stars.
@@ II.2 Numerical approach
where $\bar{W}=W/r^{2}$. The parameter space of the parity-odd RBS, depicted in Figure 1 where a typical inspiralling pattern is shown over the frequency range $w_{\text{min}}<w<1$ with $w_{\text{min}}\approx 0.6835$, has 13 marked points that serve as the typical cases to be discussed later in the paper, and the specific parameters of these points are presented in Table 1.
Figure 1: Parity-odd RBS solutions in an Arnowitt-Deser-Misner (ADM) mass $M$ vs scalar field frequency $w$ diagram. Points $1$-$13$ corresponds to the BSs to be discussed.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants Nos. 12275106 and 12235019), the Shandong Provincial Natural Science Foundation (Grant No. ZR2024QA032), and the Innovation Grogram of Shanghai Normal University (Grant No. KF202472). Some of the calculations were performed in High-performance Computing Platform at University of Jinan. YH would like to thank Professor Carlos Herdeiro for his prompt and helpful correspondence.